The son is three years older than the daughter.
In 17 years the son will be twice as old as the daughter is now.
What is the daughters present age?
Can someone please show me all the steps and what the answer is please?

Answer by Squaretimes47(1)

I am going to be using s to represent son and d to represent daughter.
s=son
d=daughter
From the first part,"(the son) (is) (three years) (older) than (the daughter)"
s = 3 + d
Then, "(In 17 years) (the son)(will be)(twice)as old as(the daughter is now.)"
17 + s = 2 d
Keep in mind that the daughter has not "aged". We know this by "the daughter is now"
You get the two equations
s=3+d
17+s=2d
and you substitute the equations
PART ONE
17+(3+d)=2d
17+3+d=2d add
20+d=2d divide
20=d
PART TWO
s=3+(20)
s=23 add
So the son is 23 now, in 17 years he will be 40, which is twice the age of the daughter, 20.